### v2

# Supervising 3D Talking Head Avatars with Analysis-by-Audio-Synthesis

Abstract
In order to be widely applicable, speech-driven 3D head avatars must articulate their lips in accordance with speech, while also conveying the appropriate emotions with dynamically changing facial expressions. The key problem is that deterministic models produce high-quality lip-sync but without rich expressions, whereas stochastic models generate diverse expressions but with lower lip-sync quality. To get the best of both, we seek a stochastic model with accurate lip-sync. To that end, we develop a new approach based on the following observation: if a method generates realistic 3D lip motions, it should be possible to infer the spoken audio from the lip motion. The inferred speech should match the original input audio, and erroneous predictions create a novel supervision signal for training 3D talking head avatars with accurate lip-sync. To demonstrate this effect, we propose THUNDER (Talking Heads Under Neural Differentiable Elocution Reconstruction), a 3D talking head avatar framework that introduces a novel supervision mechanism via differentiable sound production. First, we train a novel mesh-to-speech model that regresses audio from facial animation. Then, we incorporate this model into a diffusion-based talking avatar framework. During training, the mesh-to-speech model takes the generated animation and produces a sound that is compared to the input speech, creating a differentiable analysis-by-audio-synthesis supervision loop. Our extensive qualitative and quantitative experiments demonstrate that THUNDER significantly improves the quality of the lip-sync of talking head avatars while still allowing for generation of diverse, high-quality, expressive facial animations.
Figure 1: THUNDER introduces a new paradigm for stochastic generation of 3D talking head avatars from speech with accurate lip articulation and diverse facial expressions. Given an input audio, a 3D animation is generated. The animation is then fed into an audio synthesis (mesh-to-speech) model, which generates an output audio representation. The input and output audio representations are compared, creating a novel audio-consistency supervision loop, which we coin as analysis-by-audio-synthesis.
1 Introduction
Animating 3D faces from speech has many applications. Examples include the entertainment industry (games and movies), applications in AR/VR/XR (virtual telepresence), e-commerce and perhaps future embodied digital assistants. In order to be widely applicable, the 3D head avatars must, however, appear and act believably and feel alive. While the methods to animate 3D avatars from speech have come a long way in recent years, there is still a considerable realism gap between real humans and 3D avatars.
Seminal approaches such as VOCA [cudeiro2019capture] or Karras et al. [karras2017audio], and others [fan2022faceformer, danecek2023emote, Peng2023_EmoTalk] are deterministic regressors, that predict 3D vertices from the input audio. Despite achieving good lip-sync, these methods produce facial animations without emotion [cudeiro2019capture, fan2022faceformer] or that are unnaturally static [Peng2023_EmoTalk, danecek2023emote]. This is in part due to the lack of large-scale and expression-rich training data and in part because a deterministic regressor cannot accurately produce animations of facial motions that only weakly correlate with the audio (such as eyebrow motion, eyeblinks, etc.). Hence, in order to generate animations that appear lively and believable, a good 3D speech-driven system must model the problem in a non-deterministic manner so that it can capture the many-to-many mapping between speech and facial animation.
Consequently, recent methods model the problem of 3D speech-driven animation stochastically using relatively small datasets of high-quality 3D scans [richard2021meshtalk, xing2023codetalker, stan2023facediffuser, thambiraja20233diface]; the small training size limits diversity and expressiveness. Other methods leverage lower-quality pseudo ground-truth (GT) extracted from larger-scale video datasets [zhao2024media2face, sun2024diffposetalk]. In pursuing diversity, however, stochastic methods vary the lip motions in ways that deviate from the audio signal.
The fundamental issue is that lip motions, for the same audio, vary with expression. We seek a method that generates diverse expressions but is faithful to the audio. So we ask the question: “What does it mean to be faithful to the audio?” This leads us to our key observation; that is, we seek 3D lip motions that actually reproduce the input audio. We formalize this notion with a new form of supervision for 3D speech-driven avatars using a novel mesh-to-speech model; see Fig. 1. Inspired by recent silent-video-to-speech (SVTS) methods [choi2023lip2speechunit, kim2023lip, kim2021lipvca], we propose, to our knowledge, the first mesh-to-speech model. Our mesh-to-speech model regresses the speech audio (or a representation thereof) solely from the animated 3D face. In other words, every generated animation now also produces a sound. Our rationale is simple: the more similar the produced sound is to the input speech, the more plausible the animation. Furthermore, failure to produce the correct sound should provide a good supervision signal.
Our system consists of two stages. In the first stage, we train the mesh-to-speech model to regress audio from facial motion. In the second stage, we train a diffusion model to output 3D facial animation from speech. In this stage, we utilize the frozen mesh-to-speech model. The generated facial animation is fed into the mesh-to-speech model, effectively reproducing the spoken audio. This generated audio (or representation thereof) is then compared to the input audio, creating a self-supervised training loop; this is illustrated in Fig. 1. We refer to this as analysis-by-audio-synthesis. Our experiments show that leveraging the mesh-to-speech model results in much better lip-sync quality, while still enabling stochastic production of expressive speech-driven animation.
In summary, our contributions are: (1) A novel mesh-to-speech model that regresses sound from 3D facial animations. (2) A new form of cross-modal audio-based self-supervision via comparison of the representations of the input speech and the animation-produced output speech. (3) A 3D speech-driven animation method based on diffusion, capable of generating a variety of facial expressions while maintaining accurate lip-sync. The code and models will be made publicly available .
2 Related work
2.1 Speech-driven 3D animation
The field of speech-driven 3D facial animation has made significant progress in the last two decades [cao2005expressive, edwards2016jali, edwards2020jalicyberpunk, xu2013practical, taylor2012dynamic, cohen2001animated]. Here we focus on the recent deep-learning based line of work [pham2017end, pham2017speech, karras2017audio, taylor2017generalized, zhou2018visemenet, cudeiro2019capture, richard2021meshtalk, fan2022faceformer, xing2023codetalker, danecek2023emote, thambiraja20233diface, stan2023facediffuser, zhao2024media2face, sun2024diffposetalk, yang2024probabilistic, Peng2023_EmoTalk, nocentini2024scantalk, aneja2024facetalk, aneja2024gaussianspeech, nocentini2024emovoca, enugi2024enhancing, fan2024unitalker].
Deterministic neural methods. Karras et al. [karras2017audio] were the first to utilize deep learning by training a temporal convolutional network (TCN) to predict 3D face vertices. VOCA [cudeiro2019capture] employs a pre-trained automatic speech recognition (ASR) network to regress face vertex offsets from audio, achieving good lip-sync for multiple speaking styles. Many follow-up methods use a similar paradigm [fan2022faceformer, bala2022imitator, FaceXHuBERT_Haque_ICMI23, Peng2023_EmoTalk, danecek2023emote, peng2023selftalk, nocentini2024scantalk, nocentini2024emovoca, xu2024kmtalk, fu2024mimic, aneja2024gaussianspeech, enugi2024enhancing, fan2024unitalker].
Stochastic neural methods. The first model to approach the problem stochastically was MeshTalk [richard2021meshtalk]. MeshTalk makes use of a pretrained discretized facial motion prior along with an explicit supervision mechanism that supports generation of motions that only have a weak correlation with the audio (eyebrow motion, etc.). CodeTalker [xing2023codetalker] upgrades the FaceFormer architecture by predicting codebook classes of a pretrained VQ-VAE. Both MeshTalk and CodeTalker employ autoregressive predicition and one can therefore sample the distribution of the next token to generate variety.
Thanks to the tremendous success of Diffusion Models in the image domain [ho2020denoising, rombach2022stable] they are now widely used for 3D human body animation [tevet2023hmd, chen2023mld], as well as speech-driven animation [stan2023facediffuser, thambiraja20233diface, sun2024diffposetalk, zhao2024media2face]. The diffusion methods also employ transformer-based audio feature extractors [baevski2020wav2vec, hsu2021hubert] to condition the denoiser. FaceDiffuser [stan2023facediffuser] employs a GRU-based denoiser, 3DiFACE [thambiraja20233diface] uses a TCN-based denoiser, and DiffPoseTalk [sun2024diffposetalk], Media2Face [zhao2024media2face] and FaceTalk [aneja2024facetalk] employ a transformer decoder architecture [vaswani2017attention] passing the audio condition to the denoiser via cross-attention.
Controlling the output animation. The task of controlling or editing the output animation is of great utility in production as manually editing animations is a laborious process. Many methods use a simple one-hot encoding of training subjects, which is mapped to a “style embedding” with a learnable layer. This embedding conditions the decoder to match the speaking style of the corresponding training subject [cudeiro2019capture, fan2022faceformer, xing2023codetalker, bala2022imitator]. This paradigm has been extended to control the emotion and emotion intensity [danecek2023emote, Peng2023_EmoTalk]. DiffPoseTalk [sun2024diffposetalk] uses contrastive learning to produce a style vector from a reference animation, lifting the subject-ID conditioning limitation. Media2Face [zhao2024media2face] employs CLIP features [Radford2021clip], which can be extracted from an image, or text prompts to condition the denoising process. 3DiFACE [thambiraja20233diface] demonstrates that the output of a diffusion model can be guided by a sparse set of keyframes. In this paper, instead of focusing on controlling the animation, we present a general method to improve lip-sync quality, which can be used in conjunction with the above control mechanisms.
Datasets. Most methods [karras2017audio, cudeiro2019capture, fan2022faceformer, xing2023codetalker, bala2022imitator] have made use of high quality 3D scans synchronized with audio, such VOCASET [cudeiro2019capture], BIWI [eth_biwi_00760] or Multiface [wuu2022multiface]. While these datasets are of high quality, they are expensive to acquire and hence are limited in number of subjects, and richness of facial expressions. The dataset design is often a limiting factor, as the samples are usually sentences taken out of context and hence the subjects do not express themselves in the most natural ways. As a consequence, even the stochastic methods that leverage these datasets [xing2023codetalker, richard2021meshtalk, stan2023facediffuser], suffer from lack of natural diversity of expressions.
Thanks to the significant improvement of in-the-wild face reconstruction systems [EMOCA:CVPR:2021, MICA:ECCV2022, filntisis2022visual, zhang2023tokenface, Retsinas_2024_CVPR_Smirk] that can now be utilized off-the-shelf, recent methods have turned to using pseudo-ground-truth (pGT) reconstructions from videos [danecek2023emote, Peng2023_EmoTalk, sun2024diffposetalk, zhao2024media2face]. While the pGT does not reach the quality of 3D scans, video datasets are numerous, allowing more data to be acquired, which benefits data-hungry stochastic models.
2.2 Silent-video-to-speech (SVTS)
To the best of our knowledge, the prediction of voice from a sequence of 3D facial shapes has never been explored. However, recent years have seen tremendous progress on the task of speech audio prediction from silent videos. Early methods [cornu2015speechfromfeaturs, Oneata2021speaker, hong2021speechrec, kim2021lipvca, Michelsanti2020vocoderbased, Yadav2021speech, Vougioukas2019videodriven, um2023facetron] focus on small in-the-lab datasets with predefined scripts and limited number of speakers (GRID [cooke2016grid], TCD-TIMIT [Harte2015tcdtimit]) or script-unconstrained but single-speaker models [prajwal2020cvrplearning].
STVS [carrillo2022svts] was the first method to produce intelligible audio on large-scale in-the-wild datasets such as LRS2 and LRS3 [afouras2018lrs3] by regressing spectrograms from mouth crops. The spectrograms are converted to the final waveform with a pretrained vocoder[choi2023lip2speechunit]. Follow-up methods improve the prediction quality [kim2023lip] by adding a surrogate ASR loss or predicting features from an ASR model [choi2023lip2speechunit]
The most recent architectures are based on diffusion, producing audio often indistinguishable from real speech [choi2023diffv2s, yemini2023lipvoicer]. These, however, are not suitable for a self-supervised reconstruction loop, due to the iterative nature of diffusion models.
Hence, thanks to its simple, yet efficient feed-forward design and a relatively straightforward architecture and training, we base our mesh-to-speech model on Choi et al.[choi2023lip2speechunit].
2.3 Audio cycle consistency
Cycle consistency losses on audio have been applied before, for tasks like voice conversion [Kaneko2020CycleGAN-VC, Kaneko2020CycleGAN-VC2, Kaneko2020CycleGAN-VC3] and disentangled representation learning [ji2021evp, Chhatre2024amuse]. Many different self-supervised audio learning methods exist (see survey [liu2022audioself]). However, cross-modal cycle consistency between audio and 3D facial motion has, to our knowledge, never been proposed.
3 Preliminaries
3.1 Face model
THUNDER uses the FLAME face model [FLAME:SiggraphAsia2017], which is a statistical 3D morphable model (3DMM). FLAME provides a compact representation of facial shapes and expressions and is defined as a function:
$M(\boldsymbol{\beta},\boldsymbol{\theta},\boldsymbol{\psi})\rightarrow(\mathbf{V},\mathbf{F}),$ (1)
where the inputs are shape coefficients $\boldsymbol{\beta}\in\mathbb{R}^{\left|\boldsymbol{\beta}\right|}$, expression coefficients $\boldsymbol{\psi}\in\mathbb{R}^{\left|\boldsymbol{\psi}\right|}$ and rotation vectors for $k=4$ joints $\boldsymbol{\theta}\in\mathbb{R}^{3k+3}$. FLAME outputs a 3D mesh with vertices $\mathbf{V}\in\mathbb{R}^{n_{v}\times 3}$ and triangles $\mathbf{F}\in\mathbb{R}^{n_{f}\times 3}$. Since we do not focus on head movement, we only use the jaw joint rotation $\boldsymbol{\theta}_{jaw}$. For brevity, we refer to the expression coefficients and jaw rotation as expression parameters $\boldsymbol{x}=[\boldsymbol{\psi}|\boldsymbol{\theta}_{jaw}]$.
3.2 Speech feature extraction
Following previous work on speech-driven facial animation [fan2022faceformer, xing2023codetalker, bala2022imitator, danecek2023emote, zhao2024media2face] we use Wav2Vec2.0 [baevski2020wav2vec] as our audio feature extractor. It consists of a temporal convolution network (TCN) that extracts audio features at 50Hz and a transformer encoder that processes these features. Similarly to previous methods, we resample the TCN feature to 25Hz and feed it to the transformer to obtain the final speech feature. Formally, we define this as: $\mathcal{A}(\mathbf{w})\rightarrow\mathbf{a}^{1:T},$ where $\mathcal{A}$ denotes the Wav2Vec 2.0 network, $\mathbf{w}$ is the input waveform, and $\mathbf{a}^{1:T}\in\mathbb{R}^{T\times d_{s}}$, with $T$ denoting the number of frames at 25Hz and $d_{s}=768$ is the feature dimension.
3.3 Speaker Embedding
A speaker embedding is a numerical representation of a speaker’s unique vocal characteristics encoded into a vector. This vector captures features such as tone, pitch, speaking style, and other vocal attributes, enabling the differentiation and recognition of individual speakers. Among other applications, speaker embeddings are used in the latest SVTS models [carrillo2022svts, kim2023lip, choi2023lip2speechunit]. Following Choi et al. [choi2023lip2speechunit], we leverage an off-the-shelf speaker embedding extractor [RealTimeVoiceCloning]. We define the extractor as: $\mathcal{S}(\mathbf{w})\rightarrow\boldsymbol{s},$ with $\mathcal{S}$ and $\boldsymbol{s}$ denoting embedding extractor and the embedding vector, respectively.
3.4 Speech Units
The term speech units refers to discrete linguistic units identified through a self-supervised learning process. Like Choi et al. [choi2023lip2speechunit], we leverage speech units predicted by a pretrained HuBERT [hsu2021hubert] model and its associated feature clustering model. First, features are extracted and subsequently clustered into $n_{\boldsymbol{U}}$ units. We define speech units as one-hot class vectors $\boldsymbol{U}\in[c_{1},c_{2},\dots,c_{n_{\boldsymbol{U}}}]$ where only one of $c_{i},i=1\dots c_{n_{\boldsymbol{U}}}$ is $1$ and the rest are 00. Speech units are predicted at 50Hz.
4 Method
Figure 2: Mesh-to-speech architecture. It takes a sequence of mouth shapes as input, along with a speaker embedding feature, to produce the output speech units and spectrograms. These are used to compute a loss (top) and to produce the reconstructed audio (bottom) using a pretrained vocoder.
THUNDER is trained in two stages. First, a mesh-to-speech network is trained to regress audio from facial motion. Second, a diffusion model is trained to output 3D facial animation. The output animation is then fed to the frozen mesh-to-speech model producing output audio representations. This allows us to design a loss between the input and output representations, which we utilize in training the talking head system.
4.1 Mesh-To-Speech
Regressing audio from 3D facial motion has, to our knowledge, not been done before. To narrow down the architecture search space, we take inspiration from a recent SVTS model, namely Choi et al. [choi2023lip2speechunit].
Architecture. The architecture consists of an input feature encoder, a conformer sequence encoder and two prediction heads, a speech unit classifier, and a mel spectrogram regressor. These two outputs contain enough information that a pretrained off-the-shelf vocoder [choi2023lip2speechunit] can turn them into output audio. Since the input is not a video, but a 3D animation, we replace the original lip video ResNet encoder that Choi et al. [choi2023lip2speechunit] use with an MLP that takes 3D lip vertex coordinates as the input. We keep the rest of the architecture the same. The mesh-to-speech architecture is depicted in Fig. 2 and can be written
$\mathcal{F}(\mathbf{V}_{m\in\mathcal{M}}^{1:T},\boldsymbol{s})\rightarrow(\widehat{\boldsymbol{M}}^{1:T},\widehat{\boldsymbol{U}}^{1:T}),$ (2)
where $\mathcal{F}(.)$ denotes the mesh-to-speech network, $\mathcal{M}$ is a subset of mouth vertices. $\widehat{\boldsymbol{M}}^{1:T},\widehat{\boldsymbol{U}}^{1:T}$ denote the output sequence of mel spectrograms and speech units. Finally, $\widehat{\boldsymbol{M}}^{1:T}$ and $\widehat{\boldsymbol{U}}^{1:T}$ can be passed to an off-the-shelf vocoder to generate the output waveform $\widehat{\mathbf{w}}$.
Supervision. We follow the same supervision scheme as Choi et al. [choi2023lip2speechunit]. The loss consists of two terms. The first term is an L1 loss between the input and output mel specrograms:
$\mathcal{L}_{mel}=\|\boldsymbol{M}^{1:T}-\widehat{\boldsymbol{M}}^{1:T}\|_{1}.$ (3)
The second term is cross entropy between the predicted speech unit classification vectors and the GT speech units:
$\mathcal{L}_{unit}=\frac{1}{T}\sum_{t=1}^{T}\sum_{c=1}^{n_{\boldsymbol{U}}}\boldsymbol{U}_{c}^{t}\ \textup{log}\ p(\widehat{\boldsymbol{U}}_{c}^{t}),$ (4)
with weights $w_{mel}=10$ and $w_{unit}=1$. The final loss is:
$\mathcal{L}_{m2s}=w_{mel}\mathcal{L}_{mel}+w_{unit}\mathcal{L}_{unit}.$ (5)
4.2 Speech-Driven Facial Animation Diffusion
Figure 3: Diffusion architecture. The upper part of the figure (green) depicts the architecture of the diffusion model and the lower part (yellow) illustrates the application of the audio consistency loss. Gray boxes indicate the input to the system. Trainable components are highlighted in orange and frozen ones in blue.
Architecture. The core of THUNDER consists of a denoiser in the form of a transformer decoder [vaswani2017attention]. Similar to other contemporary 3D animation methods (e.g. Zhao et al. [zhao2024media2face] and Sun et al. [sun2024diffposetalk]), THUNDER processes a noisy sequence of expression parameters $\boldsymbol{\tilde{x}}_{(d)}^{1:T}$, where $d$ represents the diffusion timestep. These parameters serve as the query inputs to the transformer. The keys and values for the transformer are comprised of the input audio features $\mathbf{a}^{1:T}$ and a diffusion step embedding $\mathcal{T}(d)$, concatenated along the temporal dimension. The function $\mathcal{T}$ is defined as a sinusoidal timestep function similar to Ho et al. [ho2020denoising]. We do not incorporate additional animation-controlling conditions (such as style vectors [cudeiro2019capture, fan2022faceformer, xing2023codetalker], emotion vectors [danecek2023emote, Peng2023_EmoTalk] or CLIP embeddings [zhao2024media2face]) as this is not the goal of this work. However, the design can easily be augmented with additional input conditions. All inputs are projected into the transformer’s feature dimension, $f=128$, using dedicated learnable affine layers. Positional encoding within the transformer’s self-attention utilizes the ALiBi mechanism [press2021train]. The cross-attention layers use a diagonal binary mask to facilitate the attention mechanism, and no additional positional encoding. The model predicts the reconstructed expression parameters $\boldsymbol{\widehat{x}}_{(0)}^{1:T}$. The transformation can be formally described by the following equation:
$\mathcal{D}(\boldsymbol{\tilde{x}}_{(d)}^{1:T},d,\mathbf{a}^{1:T})\rightarrow\boldsymbol{\widehat{x}}_{(0)}^{1:T}.$ (6)
Here, $d\in\{1,\dots,D\}$ indicates the diffusion timestep, and $\boldsymbol{\widehat{x}}_{(0)}^{1:T}$ is the denoised prediction. Unlike some methods that predict noise, this model directly predicts the fully denoised space, which is used to compute the 3D mesh sequence: $M(\boldsymbol{\beta}=\mathbf{0},\boldsymbol{\widehat{x}}_{(0)}^{1:T})\rightarrow\widehat{\mathbf{V}}_{(0)}^{1:T}$. This approach allows the model to define losses directly on the final 3D mesh space, crucial for leveraging mesh-to-speech. See Fig. 3 for an overview.
Training. During training, a diffusion step $d$ is sampled and $\boldsymbol{\tilde{x}}_{(d)}^{1:T}$ is computed from the clean expression parameters $\boldsymbol{x}^{1:T}$ using a noise schedule that blends the original data with Gaussian noise based on predefined $\alpha$ values:
$\boldsymbol{\tilde{x}}_{(d)}^{1:T}=\sqrt{\alpha_{d}}\boldsymbol{x}^{1:T}+\sqrt{1-\alpha_{d}}\cdot\epsilon,\ \epsilon\sim\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})),$ (7)
[…]
$\mathcal{L}_{rec}=$ $\ \textup{MSE}(\mathbf{V}^{1:T},\widehat{\mathbf{V}}_{(0)}^{1:T})+\textup{MSE}(\boldsymbol{\psi}^{1:T},\boldsymbol{\widehat{\psi}}_{(0)}^{1:T})+$
$\ \textup{MSE}(\boldsymbol{\theta}_{jaw}^{1:T},{\boldsymbol{\widehat{\theta}}_{jaw}^{1:T}}_{(0)}).$
[…]
$\mathcal{L}_{vel}=$ $\ \textup{MSE}(v(\mathbf{V}^{1:T}),v(\widehat{\mathbf{V}}_{(0)}^{1:T}))+\textup{MSE}(v(\boldsymbol{\psi}^{1:T}),v(\boldsymbol{\widehat{\psi}}_{(0)}^{1:T}))+$
$\ \textup{MSE}(v(\boldsymbol{\theta}_{jaw}^{1:T}),v({\boldsymbol{\widehat{\theta}}_{jaw}^{1:T}}_{(0)})),$
where $v$ computes the velocity of any input sequence $\mathbf{u}^{1:T}$ $v(\mathbf{u}^{1:T})=\mathbf{u}^{2:T}-\mathbf{u}^{1:T-1}$. The final loss is given as:
$\mathcal{L}_{total}=w_{m2s}\mathcal{L}_{m2s}+\mathcal{L}_{rec}+\mathcal{L}_{vel},$ (8)
with $w_{mts}=1$. The model is trained using classifier-free guidance [ho2022cfg]. The audio condition is randomly dropped and replaced with a learnable vector with 20% probability.
Inference. At inference time, we initialize the first step by sampling the Gaussian noise: $\boldsymbol{\widehat{x}}_{(D)}\sim\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$. We then proceed to denoise for $D$ diffusion steps. In each step, we first employ the denoiser $\mathcal{D}(\boldsymbol{\tilde{x}}_{(d)}^{1:T},d,\mathbf{a}^{1:T})$ to obtain $\boldsymbol{\widehat{x}}_{(0)}^{1:T}$. Then, we compute $\boldsymbol{\tilde{x}}_{(d-1)}$ by adding the expected amount of noise back to obtain $\boldsymbol{\tilde{x}}_{(d-1)}$. Since THUNDER is trained with classifier-free guidance, it is possible to trade-off fidelity of the lip-sync for increased diversity by combining $s_{a}\mathcal{D}(\boldsymbol{\tilde{x}}_{(d)}^{1:T},d,\mathbf{a}^{1:T})+(1-s_{a})\mathcal{D}(\boldsymbol{\tilde{x}}_{(d)}^{1:T},d,\O)$ during the denoising process. However, we find our results are diverse enough without this and so we effectively set $s_{a}=1$.
5 Experiments
5.1 Datasets
The selection of datasets for THUNDER is challenging. On the one hand, 4D scan datasets such as VOCASET [cudeiro2019capture] provide good reconstructions but are too small to train generative models. On the other hand, in order to have a high-enough quality geometry from which speech can be predicted, we cannot rely on in-the-wild datasets, as reconstruction artifacts decrease the quality of the GT and hence our mesh-to-speech and talking head models.
Hence, we use a collection of public in-the-lab captured video datasets commonly used for lip-reading, namely TCD-TIMIT [Harte2015tcdtimit], RAVDESS [livingstone_2018_1188976] and GRID [cooke2016grid], totaling 117 subjects, uttering 42140 short sentences, totaling 3535816 frames at 25FPS (over 39 hours). These datasets provide clear imagery without in-the-wild complexity and can be reconstructed by off-the-shelf methods to a satisfactory degree. Following EMOTE [danecek2023emote], we utilize the publicly available INFERNO framework [inferno2023], which contains a SOTA face reconstruction system [MICA:ECCV2022, EMOCA:CVPR:2021, filntisis2022visual]. We split each of the datasets by subjects, 70% for training and 15% each for validation and testing.
5.2 Mesh-To-Speech
Implementation details. We initialize the conformer and prediction heads with pretrained weights from Choi et al. [choi2023lip2speechunit]. The input MLP is trained from scratch. We train the model with the Adam optimizer [Kingma2015] for 20 epochs and select the checkpoint with the lowest validation loss (reached within 5 epochs). The training batch size is 16 and the input sequence is trimmed to 125 frames.
Quantitative comparison. Since the mesh-to-speech model is the first of its kind, we compare mesh-to-speech to a recent silent-video-to-speech model [choi2023lip2speechunit] which we finetune on the videos from our training dataset for fair comparison. Note that FLAME does not model the teeth or tongue, which makes the mesh-to-speech task arguably harder than video-to-speech. Teeth and tongue, which may be visible on video, are critical for production of certain phonemes, such as labiodentals (/f/, /v/), or alveolar phonemes (/s/, /z/, /t/, /d/, /l/) [Ladefoged2011phonetics]. Consequently, we do not expect mesh-to-speech to match the performance of a video-based method.
Table 1 reports the standard video-to-speech metrics, namely Short-Time Objective Intelligibility [Taal2011stoi], Extended STOI [Jensen2016estoi] to measure the intelligibility of the generated samples, Perceptual Evaluation of Speech Quality (PESQ) [Rix2001pesq] (narrow and wide band) to measure the perceptual quality, and Word Error Rate (WER), which we measure with the SOTA ASR network Whisper[radford2023whisper]. As expected, mesh-to-speech is not as accurate as video-to-speech, but the performance is comparable, suggesting that there is sufficient audio signal in the 3D to be useful for our downstream task. Note that mesh-to-speech is, however, more accurate than Choi et al.’s original (but not the fine-tuned) model. We experiment with the following input spaces for mesh-to-speech - selection of mouth vertices $\mathbf{V}_{m\in\mathcal{M}}$ (mouth2s), all vertices $\mathbf{V}$ (exp2s), and FLAME expression parameters $\boldsymbol{x}$ (exp2s) and report the results in Tab. 1. Remarkably, the model which takes global expression vectors, performs slightly better than the selection of mouth vertices $\mathbf{V}_{m\in\mathcal{M}}$. Despite that, we choose mouth2s to be our final model, as it results in better lip-sync supervision (discussed in Sec. 5.3 and Tab. 2). Please refer to the Sup. Video for audible comparison and Sec. 7.1 in Sup. Mat. for further discussion on performance.
Modality Model: STOI $\uparrow$ ESTOI $\uparrow$ PESQ-WB $\uparrow$ PESQ-NB $\uparrow$ WER $\downarrow$
Mesh-to-speech exp2speech $\boldsymbol{x}$ 0.502 0.273 1.257 1.468 0.648
face2speech $\mathbf{V}$ 0.458 0.203 1.225 1.437 0.953
mouth2speech $\mathbf{V}_{m\in\mathcal{M}}$ 0.506 0.272 1.246 1.457 0.678
Video-to-speech Choi et al. (finetuned) 0.555 0.348 1.281 1.511 0.437
Choi et al. (orig) 0.376 0.141 1.126 1.313 1.011
Table 1: Quantitative comparison of video-to-speech and forms of mesh-to-speech (mouth2speech, face2speech and exp2speech).
5.3 Speech-Driven Facial Animation
Implementation details. THUNDER is trained using the Adam optimizer for 260 epochs with batch size 48, sequence length of 70 frames, and $D=1000$ diffusion steps with a linear noise schedule. The transformer has 8 layers with 4 heads and feature dim $f=128$.
Baselines. We compare THUNDER to other speech-driven avatar methods that are also trained on pGT and which output 3DMM parameters (as opposed to vertices). We reimplemented and trained several SOTA methods, namely FlameFormer (an adaptation of FaceFormer [fan2022faceformer]), Media2Face [zhao2024media2face] and DiffPoseTalk [sun2024diffposetalk] (see Sup. Mat. for details). Unless stated otherwise, we keep the pretrained Wav2Vec weights frozen. Models that finetune Wav2Vec are denoted with a suffix -T (T for trainable).
Evaluation protocol. For each test audio sequence, we generate S=32 outputs. Each of these is initialized with a different random noise in the case of diffusion models, or randomly sampled subject-id condition in case of FlameFormer. Then we compute the following evaluation metrics, averaged out across the S outputs (unless stated otherwise).
Lip Vertex Error. Following [richard2021meshtalk, sun2024diffposetalk, xing2023codetalker] we report LVE, which calculates the maximum L2 error of all lip vertices for each frame and then computes the average over all frames.
Dynamic Time Warping. We also report the DTW proposed by Thambiraja et al. [bala2022imitator]. First, distances between the mid-points of lower and upper lips are calculated for both the predictions and GT. The two resulting time series are then used to compute the DTW distance.
Lip Correlation Coefficients. We argue that the above metrics are not comprehensive enough as LVE only operates over a sparse set of lip vertices and does not model the temporal aspect very well as each frame is handled separately and DTW only models the lower-upper lip distance. Specifically, we compute Pearson and Concordance correlation coefficients for every vertex coordinate in the mouth-region and average the coefficients. PCC and CCC are commonly used metrics for comparing time series.
Face dynamic deviation. An important aspect of facial animation, is the expressiveness of the upper face vertices. Following CodeTalker [xing2023codetalker], we report Face Dynamics Deviation (FDD), which measures the differences between the temporal standard deviation of the pGT and the predicted vertices, averaged over a set of vertices. We report upper-FDD (FDD-U) and lip-FDD (FDD-L).
Sample diversity. One of the most important aspects of stochastic models is that they generate diverse animations for the same input. Remarkably, previous work has not analyzed this aspect [stan2023facediffuser, sun2024diffposetalk, zhao2024media2face]. We compute standard deviations of vertex distances between the pGT and the predictions for the same input at frame $t$. Formally, given a tensor of per-vertex L2 distances $\mathbf{D}\in\mathbb{R}^{N\times S\ \times T\times n_{v}}$, with $N$ being the size of the test dataset and $S=32$ is the number of sampled outputs per input. We compute the standard deviation along the sample dimension $S$. Other dimensions are subsequently averaged. We measure sample diversity for the upper face region (S-Div-U) and the lip region (S-Div-L).
Temporal diversity metric. We also report the temporal diversity, where the standard deviation is computed over the temporal dimension of $\mathbf{D}\in\mathbb{R}^{N\times S\ \times T\times n_{v}}$ (and averaged across the rest). Again, we calculate the diversity for both lip (T-Div-L) and upper face regions (T-Div-U).
5.3.1 Quantitative evaluation
Tab. 2 reports the computed metrics computed on the test set for THUNDER, baselines and ablations.
Experiment Model LVE $\downarrow$ L-CCC $\uparrow$ L-PCC $\uparrow$ DTW $\downarrow$ S-DIV-U $\uparrow$ T-DIV-U $\uparrow$ S-DIV-L $\downarrow$ T-DIV-L $\uparrow$ T-FDD-U $\downarrow$ T -FDD-L $\downarrow$
THUNDER and mesh-to-speech input space THUNDER w/o m2s 0.879 0.359 0.568 0.329 0.0419 0.044 0.21 0.254 0.0118 0.0932
THUNDER w/ face2s 0.804 0.411 0.633 0.285 0.0297 0.0409 0.128 0.237 0.0117 0.0827
THUNDER w/ exp2s 0.83 0.362 0.63 0.296 0.0404 0.0398 0.176 0.228 0.0125 0.0939
THUNDER w/ mouth2s 0.802 0.426 0.639 0.29 0.0322 0.04 0.134 0.241 0.0122 0.0806
THUNDER and trainable audio enc. THUNDER-T w/o m2s 0.723 0.428 0.623 0.266 0.02 0.0411 0.0656 0.216 0.0118 0.0811
THUNDER-T w/ mouth2s 0.709 0.445 0.66 0.256 0.021 0.039 0.0669 0.202 0.0118 0.0788
Mesh-to-speech with other methods and additional input conditions FlameFormer* 0.809 0.368 0.57 0.291 0.0271 0.0372 0.132 0.239 0.0117 0.0909
FlameFormer* w/ m2s 0.794 0.411 0.614 0.265 0.0263 0.0393 0.111 0.219 0.0107 0.0715
Media2Face* 0.96 0.308 0.531 0.363 0.0163 0.0459 0.108 0.256 0.0105 0.0899
Media2Face* w/ m2s 0.815 0.428 0.609 0.282 0.0128 0.0586 0.0371 0.233 0.0154 0.083
DiffPoseTalk* 0.68 0.464 0.612 0.267 0.025 0.0379 0.117 0.223 0.0101 0.0776
DiffPoseTalk* w/ m2s 0.651 0.469 0.653 0.268 0.0172 0.0377 0.0604 0.204 0.0117 0.0813
Table 2: Quantitative evaluation of THUNDER, other models and ablations. (1) Mesh-to-speech input space. All of the mesh-to-speech variants are applicable. THUNDER with mouth2s results in the best lip-sync performance (LVE, L-CCC, L-PCC), which is why we select it as the final THUNDER model. exp2s is applicable, too as its scores best on DTW and has good diversity (S-DIV-U), while face2s scores the worst on diversity (S-DIV-U, T-DIV-U). (2) Trainable audio encoder. THUNDER-T results in improved lip-sync metrics (LVE, L-CCC, L-PCC and DTW), but at the slight expense of diversity of outputs (mainly upper face diversity S-DIV-U and T-DIV-U) compared to THUNDER which does not train the audio encoder. (3) Mesh-To-Speech applied to other methods. The application of mesh-to-speech to our reimplementations of Media2Face* (diffusion with image conditioning), DiffPoseTalk* (diffusion with style feature conditioning) and FlameFormer (deterministic with subject ID conditioning) improves lip-sync metrics at slight expense of expression diversity. Please note that DiffPoseTalk* and Media2Face* (both with and without mesh-to-speech) have a considerably lower diversity than THUNDER. This happens due to the additional conditioning inputs which narrow down the output distribution space. This also manifests itself with better lip-sync performance.
Ablation study. We ablate our design decisions. (1) The input space for the mesh-to-speech loss, (2) the trainability of the audio encoder (Tab. 2) , and (3) the weight of the mesh-to-speech loss (see Sec. 5 in Sup. Mat. and video).
Mesh-to-speech loss with other methods. Tab. 2 shows that mesh-to-speech is applicable to other methods. We re-implement three methods that incorporate additional conditioning mechanisms and train them with and without our mesh-to-speech loss. Our new loss improves performance on the lip-sync metrics. For visual results, refer to Sup. Video. Comparison to additional methods. Tab. 7 in Sup. Mat. we also include a comparison of THUNDER and other (original) implementations recent methods.
Disentanglement effect. To test whether mesh-to-speech helps preserve the effect of good lip-sync in the presence of other editing conditions, we additionally run our reimplementations of Media2Face with conditioning inputs extracted from emotional images. See Sec. 7.2.1 in Sup. Mat.
Other datasets. To verify, whether THUNDER is dataset-agnostic, we also trained and tested the method on TFHP. We report the results in Sec. 7.2.4 in Sup. Mat..
Perceptual study. We conduct a perceptual study in which we evaluate THUNDER and the effect of mesh-to-speech. Specifically, we compare to the deterministic FlameFormer, THUNDER with and without mesh-to-speech, and p-GT. The participants are shown two videos side-by-side and asked to rate animations on a five-point Likert scale (strong/weak preference for A or B or indifferent). We ask the participants to rate three aspects of the animation: lip-sync quality, dynamism, and realism. Fig. 4 reports the results for THUNDER and Fig. 5 for THUNDER-T. For the remaining results of the study and further discussion, see Sec. 13 in Sup. Mat.
Figure 4: Perceptual study of THUNDER. We compare THUNDER-F with methods having both a trainable audio encoder (THUNDER-T w/o m2s) and frozen encoders (FlameFormer-F, THUNDER-F w/o m2s), and GT. The participants prefer THUNDER-F’s lip-sync over that of other models, but does not quite reach the fidelity of GT. Remarkably, the participants also have slight preference for THUNDER-F over GT, which suggests that the application of mesh-to-speech helps THUNDER saturate the quality of GT.
Figure 5: Perceptual study of THUNDER-T. We compare THUNDER-T with methods having both a trainable audio encoders (FlameFormer-T, THUNDER-T w/o m2s) and GT. Models with trainable backbones, achieve good lip-sync than their frozen counterparts, but THUNDER-T is still preferred over the other models.
5.3.2 Qualitative evaluation
Figure 6: Qualitative comparison. This figure shows the comparison between baselines, our model and GT for selected utterances. The method’s training dataset is listed in parantheses below.
We demonstrate the superiority of our method qualitatively. Fig. 6 shows the comparison of our method with the baselines for selected utterances. Additional qualitative evaluation, can be found in the supplemental video and PDF.
5.4 Limitations
THUNDER has the following limitations. The absence of teeth and tongue inherently creates ambiguity, making it more difficult to produce high quality audio compared to silent video-to-speech. Further, pseudo-GT is not a perfect reconstruction but has occasional artifacts or inaccuracies. Also, our current mesh-to-speech architecture is fully deterministic and hence is not completely capable of capturing the prediction ambiguities. All of these may result in suboptimal effects of the audio cycle-consistency loss. However, regardless of these shortcomings, our experiments demonstrate the benefits of our method. Employing a large-scale dataset of high quality 3D scans, a face model that models teeth and tongue, and upgrade mesh-to-speech to a stochastic model are likely to boost the benefits of the mesh-to-speech loss. Finally, the inference process of the diffusion model is currently computationally intensive. Future work should address this by employing more efficient solvers such as DPM++ [lu2022dpmpp].
6 Conclusion
In this paper, we introduced THUNDER, a 3D speech-driven avatar generation system with a novel paradigm of supervision via analysis-by-audio-synthesis. THUNDER has two key components. First, drawing inspiration from silent video-to-speech methods, we define a new mesh-to-speech task and train a model that regresses speech from facial animation. Second, we incorporated mesh-to-speech into a diffusion-based 3D speech-driven avatar system, creating the first-of-its-kind audio-based self-supervision loop that helps ensure that the produced lip animation is plausible for the spoken audio. Our extensive quantitative, perceptual, and qualitative experiments demonstrate that THUNDER achieves a significant improvement in lip animation quality for diffusion-based 3D speech-driven avatars while still producing a diverse set of facial animations. Furthermore, we have demonstrated that the application of our mesh-to-speech loss improves the lip-sync quality of other talking head avatar systems. Finally, we believe that our novel audio-based self-supervision paradigm can impact other applications of 3D head avatars, such as video-based reconstruction, automatic postprocessing of 3D facial animation, or as a quality measure of talking head avatars.
[…]
Supplementary Material
7 Additional Experiments
7.1 Mesh-to-Speech
Per-Dataset Performance Analysis. The performance of both mesh-to-speech and video-to-speech models varies dramatically depending on the dataset. Datasets with limited number of utterances and vocabulary such as RAVDESS or GRID are less challenging, and the models can regress relatively accurately, sometimes to the point of perfect intelligibility. On the more challenging TCD-TIMIT dataset, which contains more complex vocabulary, words are often missed. However, despite the relatively high word error rate (WER) on TCD-TIMIT, the produced sounds, remain very realistic, given the animation. Please see Tab. 3 for per-dataset breakdown and refer to the supplementary video for an audible comparison.
DATASET Model: STOI $\uparrow$ ESTOI $\uparrow$ PESQ-WB $\uparrow$ PESQ-NB $\uparrow$ WER $\downarrow$
GRID exp2speech $\boldsymbol{x}$ 0.533 0.290 1.298 1.552 0.553
face2speech $\mathbf{V}$ 0.482 0.211 1.261 1.513 0.783
mouth2speech $\mathbf{V}_{m\in\mathcal{M}}$ 0.538 0.295 1.288 1.539 0.559
Choi et al. (finetuned) 0.590 0.371 1.324 1.597 0.428
Choi et al. 0.426 0.158 1.137 1.345 1.202
RAVDESS exp2speech $\boldsymbol{x}$ 0.460 0.284 1.159 1.262 0.145
face2speech $\mathbf{V}$ 0.474 0.293 1.164 1.266 0.173
mouth2speech $\mathbf{V}_{m\in\mathcal{M}}$ 0.510 0.328 1.161 1.262 0.131
Choi et al. (finetuned) 0.548 0.375 1.174 1.277 0.060
Choi et al. 0.452 0.245 1.074 1.149 0.774
TCD-TIMIT exp2speech $\boldsymbol{x}$ 0.409 0.206 1.146 1.249 0.929
face2speech $\mathbf{V}$ 0.368 0.139 1.117 1.228 1.431
mouth2speech $\mathbf{V}_{m\in\mathcal{M}}$ 0.389 0.174 1.131 1.243 1.070
Choi et al. (finetuned) 0.436 0.257 1.168 1.293 0.579
Choi et al. 0.172 0.041 1.103 1.254 0.641
COMBINED exp2speech $\boldsymbol{x}$ 0.502 0.273 1.257 1.468 0.648
face2speech $\mathbf{V}$ 0.458 0.203 1.225 1.437 0.953
mouth2speech $\mathbf{V}_{m\in\mathcal{M}}$ 0.506 0.272 1.246 1.457 0.678
Choi et al. (finetuned) 0.555 0.348 1.281 1.511 0.437
Choi et al. (orig) 0.376 0.141 1.126 1.313 1.011
Table 3: Quantitative comparison of mesh-to-speech with silent-video-to-speech. While Choi et al. outperforms THUNDER in most cases, the performance is comparable. The performance of both methods depends dramatically on the test dataset. Inputs from datasets with limited vocabulary (GRID, RAVDESS) tend to produce lower word error rates than datasets with richer vocabulary (TCD-TIMIT).
7.2 Talking Heads
7.2.1 Disentanglement
Ideally, mesh-to-speech should help preserve high-quality lip-sync in the presence of other editing conditions. In other words, the lip-sync should effectively be disentangled from other conditions. To validate this, we evaluate our re-implementations of Media2Face on our test set, passing additional conditioning inputs extracted from emotional images. The conditioning images can be found in Fig. 7. For example results of Media2Face animations, please refer to the supplementary video. The metrics can be found in Tab. 4.
Name LVE $\downarrow$ LIP CCC $\uparrow$ LIP PCC $\uparrow$ DTW $\downarrow$ S-DIV-L $\downarrow$
M2F-F* orig. 0.611 0.522 0.617 0.221 0.0956
M2F-F* w/ m2s orig. 0.551 0.558 0.679 0.212 0.027
M2F-F* angry 1.6 0.307 0.526 0.914 0.151
M2F-F* w/ m2s angry 1.46 0.348 0.59 0.741 0.08
M2F-F* calm 0.91 0.369 0.537 0.297 0.103
M2F-F* w/ m2s calm 0.86 0.459 0.638 0.302 0.0289
M2F-F* disgust 0.928 0.314 0.527 0.316 0.0675
M2F-F* w/ m2s disgust 0.874 0.426 0.627 0.305 0.019
M2F-F* fearful 1.01 0.38 0.618 0.315 0.101
M2F-F* w/ m2s fearful 1.03 0.353 0.627 0.303 0.0295
M2F-F* happy 1.31 0.201 0.394 0.484 0.152
M2F-F* w/ m2s happy 1.11 0.3 0.597 0.269 0.047
M2F-F* sad 0.96 0.308 0.531 0.363 0.108
M2F-F* w/ m2s sad 0.815 0.428 0.609 0.282 0.0371
M2F-F surprised 1.04 0.391 0.576 0.494 0.0759
M2F-F w/ m2s surprised 1.06 0.366 0.602 0.362 0.0278
Table 4: Disentanglement effect. We feed different input conditions to our versions of Media2Face* models and report the lip-sync metrics for both. Orig. denotes an image condition extracted from the video that corresponds to the audio. All other rows are results for denoising with other conditions that come from 7 different emotional images. Media2Face* with THUNDER exhibits superior lip-sync in most cases.
Figure 7: Media2Face conditioning images. These images were used as the conditions for the Media2Face* disentanglement experiment in Tab. 4. The images were selected out from the RAVDESS test set. Top row from left to right: happy, angry, sad, disgusted. Bottom row: fearful, surprised, calm.
7.2.2 Extended Ablation and Sensitivity Analysis
Name LVE $\downarrow$ L-CCC $\uparrow$ L-PCC $\uparrow$ DTW $\downarrow$ S-DIV-U $\uparrow$ T-DIV-U $\uparrow$ S-DIV-L $\downarrow$ T-DIV-L $\uparrow$ FDD-U $\downarrow$ FDD-L $\downarrow$
THUNDER w/o m2s 0.879 0.359 0.568 0.329 0.0419 0.044 0.21 0.254 0.0118 0.0932
THUNDER w/ face2s 0.804 0.411 0.633 0.285 0.0297 0.0409 0.128 0.237 0.0117 0.0827
THUNDER w/ exp2s 0.83 0.362 0.63 0.296 0.0404 0.0398 0.176 0.228 0.0125 0.0939
THUNDER w/ mouth2s 0.802 0.426 0.639 0.29 0.0322 0.04 0.134 0.241 0.0122 0.0806
THUNDER-F w/o m2s 0.879 0.359 0.568 0.329 0.0419 0.044 0.21 0.254 0.0118 0.0932
THUNDER-F $w_{m2s}=0.1$ 0.887 0.375 0.599 0.309 0.0422 0.0445 0.211 0.258 0.0128 0.0936
THUNDER-F $w_{m2s}=0.5$ 0.805 0.342 0.574 0.294 0.0387 0.0383 0.164 0.225 0.0131 0.103
THUNDER-F $w_{m2s}=1.0$ 0.802 0.426 0.639 0.29 0.0322 0.04 0.134 0.241 0.0122 0.0806
THUNDER-F $w_{m2s}=5.0$ 0.858 0.32 0.572 0.313 0.0257 0.0364 0.159 0.23 0.0143 0.105
THUNDER-F $w_{m2s}=10.0$ 0.886 0.311 0.585 0.31 0.021 0.035 0.146 0.219 0.0155 0.105
THUNDER-T w/o m2s 0.723 0.428 0.623 0.266 0.02 0.0411 0.0656 0.216 0.0118 0.0811
THUNDER-T $w_{m2s}=0.1$ 0.79 0.417 0.63 0.252 0.0279 0.0431 0.0964 0.221 0.0117 0.0729
THUNDER-T $w_{m2s}=0.5$ 0.693 0.435 0.641 0.263 0.0221 0.0391 0.0662 0.201 0.0118 0.0782
THUNDER-T $w_{m2s}=1.0$ 0.709 0.445 0.66 0.256 0.021 0.039 0.0669 0.202 0.0118 0.0788
THUNDER-T $w_{m2s}=5.0$ 0.844 0.343 0.614 0.262 0.0213 0.0351 0.0997 0.203 0.0139 0.0908
THUNDER-T $w_{m2s}=10.0$ 0.737 0.408 0.621 0.301 0.0207 0.0344 0.0486 0.199 0.0147 0.089
Table 5: Ablation study and sensitivity analysis. Here we analyze the effect of the mesh-to-speech loss on the talking head avatar training. The top section compares models supervised with mouth-mesh-to-speech (mouth2s), full-mesh-to-speech (face2s) and flame-to-speech (exp2s). While it is already present in the main paper text, we include it here for completeness. The mid and bottom sections analyze the effect of weight $w_{mts}$, with either frozen (-F) or trainable (-T) Wav2Vec2. THUNDER-T models exhibit superior lip-sync performance but it comes at the expense of reduced diversity. Furthermore, for both THUNDER-F and THUNDER-T models, increasing $w_{mts}$, while beneficial for lip animation quality, comes at an increasing expense of of diversity. Please refer to the Sup. Video for qualitative comparison.
Tab. 5 shows the complete ablation of all THUNDER components. We give a comprehensive evaluation of our design decisions. (1) The input space for the mesh-to-speech loss, (2) the weight of the mesh-to-speech loss, and (3) the trainability of the input audio encoder.
On mesh-to-speech input space. While all three modalities are comparable, mouth2s has scores best on the lip-sync metrics, likely thanks to the localized effect on the mouth. Exp2s supervision performs slightly worse on lip-sync but results in higher sample diversity. Finally, applying supervision through the whole face to speech (face2s), while achieving comparable lip-sync scores, scores worse on diversity, specifically S-DIV-U.
On mesh-to-speech strength. Applying mesh-to-speech loss, even weakly, results in improved lip-sync metrics (LVE, CCC, PCC and DTW). The stronger the loss, the more the lip-sync metrics improve. However, increasing weights of the mesh-to-speech loss come at the expense of generation diversity (metrics S-DIV, T-DIV).
On finetuning the audio encoder. Nearly all recent methods finetune the input audio encoder or else they suffer from considerably impaired lip-sync. This results in a certain amount of overfitting to the training subjects’ voices, which manifests as lower diversity scores (S-DIV-U) compared to the models with frozen Wav2Vec (THUNDER-F). In essence, the model’s with trainable audio encoders become less stochastic and more deterministic. Remarkably, the application of the mesh-to-speech loss alleviates the necessity for finetuning Wav2Vec, producing significantly better lip-sync even without finetuning Wav2Vec (THUNDER-T), dramatically reducing the number of training parameters. Applying mesh-to-speech along with finetuning Wav2Vec results in further improvement of lip-sync metrics.
Please refer to the supplementary video for visual examples of all the phenomena described above.
7.2.3 Perceptual Study
Figure 8: THUNDER-F perceptual study - lip-sync. The results of THUNDER-F against other methods. We compare THUNDER-F against methods with both trainable audio encoder ( FlameFormer-T, THUNDER-T w/o mesh-to-speech) and frozen encoders (FlameFormer-F, THUNDER-F w/o mesh-to-speech). THUNDER-F outperforms other methods on lip-sync, including those with the trainable audio-encoder (denoted -T).
Figure 9: THUNDER-F perceptual study - dynamism The participants found THUNDER-F considerably more dynamic that the deterministic FlameFormer and also more dynamic than THUNDER THUNDER-F w/o mesh-to-speech. However, ground truth is still rated more dynamic.
Figure 10: THUNDER-F perceptual study - realism The participants found THUNDER-F considerably more dynamic that the deterministic FlameFormer and also more dynamic than THUNDER-F w/o mesh-to-speech. THUNDER-T w/o mesh-to-speech is rated about as realistic as THUNDER. Remarkably, the participants have preferred THUNDER’s generation over ground truth in terms of realism..
Figure 11: THUNDER-T complete perceptual study. The results of THUNDER-T against other methods. THUNDER-T outperforms its counterpart with mesh-to-speech and the deterministic FlameFormer on lip-sync quality. Additionally, the participatns have a slight preference for THUNDER-T over ground truth. This suggests, that while THUNDER-T exhibits a degree of dynamis, the GT potential has not yet been fully exhausted.
Figure 12: THUNDER-T perceptual study - dynamism. The participants rate THUNDER-T to be more dynamic than THUNDER-T without mesh-to-speech and also more dynamic than FlameFormer-T, THUNDER-T w/o mesh-to-speech and even Ground Truth. This suggest that THUNDER-T’s realism exhausted the realism of the dataset.
Figure 13: THUNDER-T perceptual study - realism. The participants prefer the degree of realism of THUNDER-T’s outputs over FlameFormer-T
Here we provide the rest of our comprehensive perceptual evaluation of THUNDER. We run a perceptual study on Amazon Mechanical Turk. The participants are shown two videos side-by-side. One corresponding to THUNDER and one to a baseline. The left-right order is randomized. The participants must finish watching both videos before being allowed to rate the videos on a five-point Likert scale (strong preference for left, weak preference for left, indifference, weak preference for right and strong preference for right). We ask the participants to rate three aspects of the animation - lip-sync, dynamism and realism. The exact task description can be seen in the study template in Fig. 14. The participants are shown 20 comparisons, generated from test audios, which where randomly selected from the RAVDESS and TCD-TIMIT test sets. In addition to that, we repeat the first 4 comparisons at the end of the study and discard the first 4 responses. This gives the participants a few examples such that they can adjust to the task. Additionally, we include 4 catch trials, where one animation is selected from Ground Truth and the other animation is clearly wrong. We discard the participants that are wrong on more than one catch trial. The complete results of the perceptual study for THUNDER for lip-sync, dynamism and realism can be found in Figures 10, 10 and 10 respectively. The results for THUNDER-T are given in Figures 13, 13 and 13. The results validate our qualitative and quantiative findings - THUNDER models exhibit better lip-sync, realism and dynamism when compared to baselines.
Figure 14: The perceptual study web template.
7.2.4 Generalization to other datasets
To verify whether the performance of THUNDER and mesh-to-speech translates to other types of data, we run additional experimetns on TFHP (the DiffPoseTalk dataset [sun2024diffposetalk]. TFHP was reconstructed from YouTube videos The dataset contains unscripted audios, longer sequences, and the reconstructions were obtained using MICA [MICA:ECCV2022] and SPECTRE [filntisis2022visual] and look qualitatively very different compared to our EMICA reconstructions. Can THUNDER work also in this setting?
We train both mesh-to-speech and THUNDER-T on TFHP, along with training DiffPoseTalk with the official DiffPoseTalk codebase. Please note, that in order to equalize the input conditions, we trained DiffPoseTalk without identity condioning. During evaluation, we do not use the style-condition but evaluate using the audio condition only (for both THUNDER models and DiffPoseTalk). The results are reported in Tab. 6.
Name LVE $\downarrow$ L-CCC $\uparrow$ L-PCC $\uparrow$ DTW $\downarrow$ S-DIV-U $\uparrow$ T-DIV-U $\uparrow$ S-DIV-L $\downarrow$ T-DIV-L $\uparrow$ T-FDD-U $\downarrow$
DiffPoseTalk (w/o style cond.) 1.24 0.341 0.508 0.25 0.0388 0.0395 0.235 0.318 0.0111
THUNDER-T w/o m2s 1.1 0.332 0.518 0.226 0.0208 0.0293 0.122 0.277 0.0068
THUNDER-T 1.09 0.351 0.545 0.214 0.0255 0.0322 0.103 0.265 0.00777
Table 6: Experiment on TFHP. We compare THUNDER-T, THUNDER w/o mesh-to-speech and DiffPoseTalk (all trained on TFHP). Similarly to our other experiments, THUNDER-T results in superior lip-sync metrics, while trading off upper-face diversity.
7.2.5 Comparison to other methods.
Comparing different speech-driven animation methods is a difficult task, especially if trained on a different dataset and especially if the data was acquired in different ways (4D scans, pseudo-GT of different reconstruction methods, etc.). Having a fair apples-to-apples comparisons stemming from measurements done against the (pseudo-) ground truth is therefore difficult. Regardless, for the sake of completeness, we compare THUNDER to the official releases of recent methods that were also trained on pseudo-GT, namely EMOTE [danecek2023emote] (a deterministic SOTA) and DiffPoseTalk [sun2024diffposetalk] (stochastic SOTA). EMOTE is a FaceFormer-inspired architecture with additional input conditions for emotions and intensity, trained with a content-emotion disentanglement mechanism. It was trained on pseudo-GT reconstructions of MEAD. The pGT was acquired with the same INFERNO tracker, making the resulting metrics comparable. The macro-architecture of DiffPoseTalk’s diffusion is very similar to THUNDER and apart from the additional style conditioning input, it differs only in small details (such as noise schedule, number of denoising steps, etc.). The most important difference is that it was trained on a different dataset, and the reconstructions were produced by a combination of SPECTRE [filntisis2022visual] and MICA [MICA:ECCV2022] and hence look qualitatively very different. As such, the DiffPoseTalk metrics likely dominated by the qualitative difference, and are only listed for completeness. Please note that for this experiment, DiffPoseTalk was run without the style condition (to match the setting of THUNDER). Finally, we include a FlameFormer trained on our data, and THUNDER trained without mesh-to-speech. The results are reported in Tab. 7.
Name LVE $\downarrow$ L-CCC $\uparrow$ L-PCC $\uparrow$ DTW $\downarrow$ S-DIV-U $\uparrow$ T-DIV-U $\uparrow$ S-DIV-L $\downarrow$ T-DIV-L $\uparrow$ FDD-U $\downarrow$ FDD-L $\downarrow$
FlameFormer* 0.809 0.368 0.57 0.291 0.0271 0.0372 0.132 0.239 0.0117 0.0909
EMOTE 1.06 0.221 0.442 0.466 0.0563 0.0401 0.248 0.24 0.0141 0.104
DiffPoseTalk 1.15 0.255 0.38 0.316 0.0345 0.0479 0.226 0.308 0.0141 0.102
THUNDER w/o m2s 0.879 0.359 0.568 0.329 0.0419 0.044 0.21 0.254 0.0118 0.0932
THUNDER 0.802 0.426 0.639 0.29 0.0322 0.04 0.134 0.241 0.0122 0.0806
Table 7: Quantitative comparison of THUNDER with other speech-driven avatar methods. Asterisk* indicates we have re-implemented the baseline and trained it on our dataset. (1) On lip-sync. THUNDER outperforms all baselines on lip-sync metrics (LVE, CCC, PCC, DTW). (2) On expression diversity. DiffPoseTalk exhibits higher temporal face diversity T-DIV-U and L-DIV-U, likely due to the training data reconstructed from videos in-the-wild, which contains more dynamic motions than our in-the-lab reconstructions. EMOTE exhibits high sample upper face diversity S-DIV-U and expression diversity S-DIV-EXP which is due to the random sampling of speaking/emotion/intensity styles, which while high, may not be fitting to the particular input audio. THUNDER w/o m2s and THUNDER follow next in S-DIV-U. DiffPoseTalk (the official model) performs comparatively low on lip-sync and well on both temporal and sample diversity, but the result is likely skewed by the fact that it was trained on different dataset with different types of reconstruction.
7.2.6 Qualitative Results
Diversity Input speech may originate from many different expression and even emotions. A well-trained stochastic model should be able to account for that and generate multiple plausible animations, possibly with different facial expressions. Fig. 15 demonstrates THUNDER’s ability to do so.
Figure 15: Diversity of outputs. THUNDER is capable of generating multiple plausible animations per audio. Each row of this figure contains 6 different generations of the same frame denoised from different initial noise samples (the audio condition remains the same for all).
Supplementary Video. Our video contains exhaustive set of results which demonstrate the benefits of mesh-to-speech and analysis-by-audio-synthesis. The reader is encouraged to watch the video for our qualitative evaluation.
8 Architecture of Baselines
Figure 16: Architecture of re-implemented baselines. (a) Our version of Media2Face [zhao2024media2face] utilized the same architecture as THUNDER but takes an extra ImabeBind feature on the input. (c) Our re-implementation of DiffPoseTalk [sun2024diffposetalk] is also based on THUNDER and takes an extra style feature from a pretrained style encoder. (c) FlameFormer is an adapted FaceFormer [fan2022faceformer]. FlameFormer predicts FLAME expression instead of full vertex space. Furthermore, FlameFormer adapts a non-autoregressive BERT-like prediction mechanism (akin to EMOTE [danecek2023emote]).
Fig. 16 shows the architecture of the baselines we reimplemented in this paper.
Media2Face. Our reimplementation of Media2Face is built on our diffusion architecture. In training, we provide additional input of ImageBind[imagebind_Girdhar2023] features extracted from images of the corresponding videos of the training set. Like the audio condition, this condition is dropped 20% of the time.
DiffPoseTalk. Our reimplementation of DiffPoseTalk is also built on our diffusion architecture. It is trained in two stages. In the first stage we train the contrastive style encoder as proposed by the authors [sun2024diffposetalk]. In training of DiffPoseTalk, we provide the additional style feature on the input. The style feature is extracted from the GT animation sequence. Like the audio condition, the style condition is dropped 20% of the time.
FlameFormer. Our FlameFormer architecture is based on FaceFormer. Like FaceFormer, FlameFormer is a transformer-based deterministic speech-driven animation network that takes addition categorical style condition on the input. Instead of predicting the full vertex space like the original FaceFormer does, FlameFormer predicts FLAME 3DMM coefficients $\boldsymbol{x}$. Additionally, instead of of using the autoregressive loop with a transformer decoder, FlameFormer’s decoder is a transformer encoder, which eliminates the need for the autoregressive formulation. This does not degrade performance and makes the model more efficient, as was already shown by Daněček et al. [danecek2023emote], which have proposed the FlameFormer baseline in their paper. Furthermore, we extend FlameFormer’s conditioning inputs - the subject identity (one-hot vector) is complemented with a one-hot vector for 8 basic emotions and three intensities.
9 Data acquisition
In this section we describe the rationale behind our choice of methodology to reconstruct the 3D faces from videos.
9.1 Choice of methodology
Acquiring high quality 4D scans of sufficient scale and richness is costly, time-consuming and requires specialized hardware. Hence, recent speech-driven animation works [sun2024diffposetalk, danecek2023emote, zhao2024media2face, yang2024probabilistic] have turned to pseudo-GT recovered from videos. This choice comes with an important decision - which face reconstruction methodolgy to use. There are several options but two of the most viable options are:
1) Optimization-based 3DDM fitting with analysis-by-image-synthesis [VetterBlanz1998, Blanz2002, Blanz2003]. While this options is viable and would produce good results, this optimization-based process is rather inefficient when deployed to many hours of video.
2) Off-the-shelf face 3DMM regressors. Due to the computational intensity of optimization-based fitting and thanks to the recent advance in deep-learning based in-the-wild face reconstruction methods [Sanyal2019_RingNet, Feng2021_DECA, EMOCA:CVPR:2021, MICA:ECCV2022, filntisis2022visual, zhang2023tokenface, Retsinas_2024_CVPR_Smirk, Tewari2017, Tewari2018, Tewari2019_FML, Deng2019]. While this is still an active area of research, and no method produces results that could be considered as accurate as 4D scans, recent years have brought enough advancement so that they can be employed for speech-driven animation dataset construction.
9.2 Discussion of FLAME regressors
Since THUNDER utilizes FLAME [FLAME:SiggraphAsia2017], our discussion will focus on FLAME regressors (trackers) only. There are several systems that have been used for talking head avatar research. No tracker Each tracker produces different kinds of errors but they were used in prior works on 3D talking head avatars. Here we briefly discuss some of the most applicable candidates:
  • DECA [Feng2021_DECA] is the first FLAME-based regressor trained with the self-supervised ”analysis-by-image-synthesis” loop. However, it is limited in terms of richness of expressions and quality of lip animations.
  • EMOCA [EMOCA:CVPR:2021] is a follow-up of DECA, capable of reconstruct rich emotions thanks to its emotion-consistency loss. However it can exaggarate expressions and the lip-sync is not good enough.
  • SPECTRE [filntisis2022visual] applies a paradigm similar to EMOCA but leverages a lip-reading consistency loss instead. It produces good lip-readable animations, but unfortunately often exaggarates, is not capable of reconstructing rich emotional expressions and the disentanglement between expression and identity is rather poor.
  • EMICA [inferno2023] is a combination of the DECA, EMOCA and SPECTRE putting together the benefits of them all to achieve high quality emotions and lip-sync. Additionally, EMICA makes use of MICA [MICA:ECCV2022] for a consistent identity prediction. The system is trained jointly and publicly available.
  • SMIRK [Retsinas_2024_CVPR_Smirk] SMIRK is a recent reconstruction method which uses a neural renderer and in-the-loop neural data augmentation to produce high quality reconstructions.
9.3 Previous works
As a result of the tremendous progress of in-the-wild face reconstruction, some of the recent SOTA speech-driven animation have already turned to pseudo-GT as the primary source of data.
  • EMOTE [danecek2023emote] employed the EMICA model to reconstruct the MEAD [kaisiyuan2020mead] dataset.
  • Yang et al. [yang2024probabilistic] utilized both DECA and SPECTRE to reconstruct LRS3 [afouras2018lrs3].
  • DiffPoseTalk employed MICA and SPECTRE separately and then combined the shape predictions of MICA and expression predictions of SPECTRE to reconstruct TFHP.
  • EMOCA exaggarates expressions.
  • DECA’s lip-sync and expressions are very limited.
  • We opt for EMICA as it provides a good compromise - rich emotions, good lip animations, consistent shape identity and temporal consistency. That said, it is not artifact-free (mouth wide open correlates with raised eyebrows, etc.) but it is sufficient to use for our research.
Following EMOTE, we opt to use EMICA in this work as well.